OpenStudy (anonymous):

A goat is tethered (tied up) to a stake in the ground with a 5 m rope. The goat can graze to the full length of the rope a full 360 around the stake. How much area does the goat have in which to graze?

jimthompson5910 (jim_thompson5910):

Use A = pi*r^2 where pi = 3.14 (approximately) r = 5

OpenStudy (anonymous):

what is the area????

OpenStudy (anonymous):

im still confuse.

jimthompson5910 (jim_thompson5910):

A = pi*r^2 A = 3.14*5^2 A = ???

OpenStudy (anonymous):

is it 78.5 ?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thank you very much !!!!! :) i guess i will pass the test tommorow